# Descrip3D: Enhancing Large Language Model-based 3D Scene Understanding with Object-Level Text Descriptions

### v1

@@ 5 Conclusion and Future Work
Reference
  • [1] Panos Achlioptas, Ahmed Abdelreheem, Fei Xia, Mohamed Elhoseiny, and Leonidas Guibas. Referit3d: Neural listeners for fine-grained 3d object identification in real-world scenes. In European conference on computer vision, pages 422–440. Springer, 2020.
  • [2] Daichi Azuma, Taiki Miyanishi, Shuhei Kurita, and Motoaki Kawanabe. Scanqa: 3d question answering for spatial scene understanding. In proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 19129–19139, 2022.
  • [3] Daigang Cai, Lichen Zhao, Jing Zhang, Lu Sheng, and Dong Xu. 3djcg: A unified framework for joint dense captioning and visual grounding on 3d point clouds. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 16464–16473, 2022.
  • [4] Dave Zhenyu Chen, Angel X Chang, and Matthias Nießner. Scanrefer: 3d object localization in rgb-d scans using natural language. In European conference on computer vision, pages 202–221. Springer, 2020.
  • [5] Dave Zhenyu Chen, Qirui Wu, Matthias Nießner, and Angel X. Chang. D3net: A speaker-listener architecture for semi-supervised dense captioning and visual grounding in rgb-d scans, 2021.
  • [6] Jiaming Chen, Weixin Luo, Xiaolin Wei, Lin Ma, and Wei Zhang. Ham: Hierarchical attention model with high performance for 3d visual grounding. arXiv preprint arXiv:2210.12513, 2(3), 2022.
  • [7] Shizhe Chen, Pierre-Louis Guhur, Makarand Tapaswi, Cordelia Schmid, and Ivan Laptev. Language conditioned spatial relation reasoning for 3d object grounding. Advances in neural information processing systems, 35:20522–20535, 2022.
  • [8] Sijin Chen, Xin Chen, Chi Zhang, Mingsheng Li, Gang Yu, Hao Fei, Hongyuan Zhu, Jiayuan Fan, and Tao Chen. Ll3da: Visual interactive instruction tuning for omni-3d understanding reasoning and planning. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 26428–26438, 2024.
  • [9] Sijin Chen, Hongyuan Zhu, Mingsheng Li, Xin Chen, Peng Guo, Yinjie Lei, Gang Yu, Taihao Li, and Tao Chen. Vote2cap-detr++: Decoupling localization and describing for end-to-end 3d dense captioning. IEEE Transactions on Pattern Analysis and Machine Intelligence, 46(11):7331–7347, 2024.
  • [10] Yilun Chen, Shuai Yang, Haifeng Huang, Tai Wang, Runsen Xu, Ruiyuan Lyu, Dahua Lin, and Jiangmiao Pang. Grounded 3d-llm with referent tokens. arXiv preprint arXiv:2405.10370, 2024.
  • [11] Zhenyu Chen, Ali Gholami, Matthias Nießner, and Angel X Chang. Scan2cap: Context-aware dense captioning in rgb-d scans. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 3193–3203, 2021.
  • [12] Wei-Lin Chiang, Zhuohan Li, Ziqing Lin, Ying Sheng, Zhanghao Wu, Hao Zhang, Lianmin Zheng, Siyuan Zhuang, Yonghao Zhuang, Joseph E Gonzalez, et al. Vicuna: An open-source chatbot impressing gpt-4 with 90%* chatgpt quality. See https://vicuna. lmsys. org (accessed 14 April 2023), 2(3):6, 2023.
  • [13] Angela Dai, Angel X Chang, Manolis Savva, Maciej Halber, Thomas Funkhouser, and Matthias Nießner. Scannet: Richly-annotated 3d reconstructions of indoor scenes. In Proceedings of the IEEE conference on computer vision and pattern recognition, pages 5828–5839, 2017.
  • [14] Rao Fu, Jingyu Liu, Xilun Chen, Yixin Nie, and Wenhan Xiong. Scene-llm: Extending language model for 3d visual understanding and reasoning. arXiv preprint arXiv:2403.11401, 2024.
  • [15] Qiao Gu, Ali Kuwajerwala, Sacha Morin, Krishna Murthy Jatavallabhula, Bipasha Sen, Aditya Agarwal, Corban Rivera, William Paul, Kirsty Ellis, Rama Chellappa, et al. Conceptgraphs: Open-vocabulary 3d scene graphs for perception and planning. In 2024 IEEE International Conference on Robotics and Automation (ICRA), pages 5021–5028. IEEE, 2024.
  • [16] Ziyu Guo, Renrui Zhang, Xiangyang Zhu, Yiwen Tang, Xianzheng Ma, Jiaming Han, Kexin Chen, Peng Gao, Xianzhi Li, Hongsheng Li, et al. Point-bind & point-llm: Aligning point cloud with multi-modality for 3d understanding, generation, and instruction following. arXiv preprint arXiv:2309.00615, 2023.
  • [17] Muhammad Usman Hadi, Rizwan Qureshi, Abbas Shah, Muhammad Irfan, Anas Zafar, Muhammad Bilal Shaikh, Naveed Akhtar, Jia Wu, Seyedali Mirjalili, et al. A survey on large language models: Applications, challenges, limitations, and practical usage. Authorea Preprints, 2023.
  • [18] Jiaming Han, Renrui Zhang, Wenqi Shao, Peng Gao, Peng Xu, Han Xiao, Kaipeng Zhang, Chris Liu, Song Wen, Ziyu Guo, et al. Imagebind-llm: Multi-modality instruction tuning. arXiv preprint arXiv:2309.03905, 2023.
  • [19] Deepti Hegde, Jeya Maria Jose Valanarasu, and Vishal Patel. Clip goes 3d: Leveraging prompt tuning for language grounded 3d recognition. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 2028–2038, 2023.
  • [20] Yining Hong, Haoyu Zhen, Peihao Chen, Shuhong Zheng, Yilun Du, Zhenfang Chen, and Chuang Gan. 3d-llm: Injecting the 3d world into large language models. Advances in Neural Information Processing Systems, 36:20482–20494, 2023.
  • [21] Edward J Hu, Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, Weizhu Chen, et al. Lora: Low-rank adaptation of large language models. ICLR, 1(2):3, 2022.
  • [22] Haifeng Huang, Yilun Chen, Zehan Wang, Rongjie Huang, Runsen Xu, Tai Wang, Luping Liu, Xize Cheng, Yang Zhao, Jiangmiao Pang, et al. Chat-scene: Bridging 3d scene and large language models with object identifiers. Advances in Neural Information Processing Systems, 37:113991–114017, 2024.
  • [23] Jiangyong Huang, Silong Yong, Xiaojian Ma, Xiongkun Linghu, Puhao Li, Yan Wang, Qing Li, Song-Chun Zhu, Baoxiong Jia, and Siyuan Huang. An embodied generalist agent in 3d world. arXiv preprint arXiv:2311.12871, 2023.
  • [24] Pin-Hao Huang, Han-Hung Lee, Hwann-Tzong Chen, and Tyng-Luh Liu. Text-guided graph neural networks for referring 3d instance segmentation. In Proceedings of the AAAI Conference on Artificial Intelligence, volume 35, pages 1610–1618, 2021.
  • [25] Shijia Huang, Yilun Chen, Jiaya Jia, and Liwei Wang. Multi-view transformer for 3d visual grounding. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 15524–15533, 2022.
  • [26] Ayush Jain, Nikolaos Gkanatsios, Ishita Mediratta, and Katerina Fragkiadaki. Bottom up top down detection transformers for language grounding in images and point clouds. In European Conference on Computer Vision, pages 417–433. Springer, 2022.
  • [27] Zhao Jin, Munawar Hayat, Yuwei Yang, Yulan Guo, and Yinjie Lei. Context-aware alignment and mutual masking for 3d-language pre-training. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 10984–10994, 2023.
  • [28] Xinyi Li, Sai Wang, Siqi Zeng, Yu Wu, and Yi Yang. A survey on llm-based multi-agent systems: workflow, infrastructure, and challenges. Vicinagearth, 1(1):9, 2024.
  • [29] Zechuan Li, Hongshan Yu, Yihao Ding, Yan Li, Yong He, and Naveed Akhtar. Embodied intelligence for 3d understanding: A survey on 3d scene question answering. arXiv preprint arXiv:2502.00342, 2025.
  • [30] Daizong Liu, Yang Liu, Wencan Huang, and Wei Hu. A survey on text-guided 3d visual grounding: Elements, recent advances, and future directions. arXiv preprint arXiv:2406.05785, 2024.
  • [31] Haotian Liu, Chunyuan Li, Qingyang Wu, and Yong Jae Lee. Visual instruction tuning. In NeurIPS, 2023.
  • [32] Junyu Luo, Jiahui Fu, Xianghao Kong, Chen Gao, Haibing Ren, Hao Shen, Huaxia Xia, and Si Liu. 3d-sps: Single-stage 3d visual grounding via referred point progressive selection. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 16454–16463, 2022.
  • [33] Xianzheng Ma, Yash Bhalgat, Brandon Smart, Shuai Chen, Xinghui Li, Jian Ding, Jindong Gu, Dave Zhenyu Chen, Songyou Peng, Jia-Wang Bian, et al. When llms step into the 3d world: A survey and meta-analysis of 3d tasks via multi-modal large language models. arXiv preprint arXiv:2405.10255, 2024.
  • [34] Xiaojian Ma, Silong Yong, Zilong Zheng, Qing Li, Yitao Liang, Song-Chun Zhu, and Siyuan Huang. Sqa3d: Situated question answering in 3d scenes. arXiv preprint arXiv:2210.07474, 2022.
  • [35] Shervin Minaee, Tomas Mikolov, Narjes Nikzad, Meysam Chenaghlu, Richard Socher, Xavier Amatriain, and Jianfeng Gao. Large language models: A survey. arXiv preprint arXiv:2402.06196, 2024.
  • [36] Maxime Oquab, Timothée Darcet, Théo Moutakanni, Huy Vo, Marc Szafraniec, Vasil Khalidov, Pierre Fernandez, Daniel Haziza, Francisco Massa, Alaaeldin El-Nouby, et al. Dinov2: Learning robust visual features without supervision. arXiv preprint arXiv:2304.07193, 2023.
  • [37] Zhangyang Qi, Zhixiong Zhang, Ye Fang, Jiaqi Wang, and Hengshuang Zhao. Gpt4scene: Understand 3d scenes from videos with vision-language models. arXiv preprint arXiv:2501.01428, 2025.
  • [38] Nils Reimers and Iryna Gurevych. Sentence-bert: Sentence embeddings using siamese bert-networks. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing. Association for Computational Linguistics, 11 2019.
  • [39] Jonas Schult, Francis Engelmann, Alexander Hermans, Or Litany, Siyu Tang, and Bastian Leibe. Mask3d: Mask transformer for 3d semantic instance segmentation. arXiv preprint arXiv:2210.03105, 2022.
  • [40] Kaitao Song, Xu Tan, Tao Qin, Jianfeng Lu, and Tie-Yan Liu. Mpnet: Masked and permuted pre-training for language understanding. Advances in neural information processing systems, 33:16857–16867, 2020.
  • [41] Ozan Unal, Christos Sakaridis, Suman Saha, Fisher Yu, and Luc Van Gool. Three ways to improve verbo-visual fusion for dense 3d visual grounding. arXiv preprint arXiv:2309.04561, 2:15, 2023.
  • [42] Yun-Cheng Wang, Jintang Xue, Chengwei Wei, and C-C Jay Kuo. An overview on generative ai at scale with edge–cloud computing. IEEE Open Journal of the Communications Society, 4:2952–2971, 2023.
  • [43] Zehan Wang, Haifeng Huang, Yang Zhao, Linjun Li, Xize Cheng, Yichen Zhu, Aoxiong Yin, and Zhou Zhao. 3drp-net: 3d relative position-aware network for 3d visual grounding. arXiv preprint arXiv:2307.13363, 2023.
  • [44] Zehan Wang, Haifeng Huang, Yang Zhao, Ziang Zhang, and Zhou Zhao. Chat-3d: Data-efficiently tuning large language model for universal dialogue of 3d scenes. arXiv preprint arXiv:2308.08769, 2023.
  • [45] Abdelrhman Werby, Chenguang Huang, Martin Büchner, Abhinav Valada, and Wolfram Burgard. Hierarchical open-vocabulary 3d scene graphs for language-grounded robot navigation. In First Workshop on Vision-Language Models for Navigation and Manipulation at ICRA 2024, 2024.
  • [46] Tung-Yu Wu, Sheng-Yu Huang, and Yu-Chiang Frank Wang. Dora: 3d visual grounding with order-aware referring. CoRR, 2024.
  • [47] Yanmin Wu, Xinhua Cheng, Renrui Zhang, Zesen Cheng, and Jian Zhang. Eda: Explicit text-decoupling and dense alignment for 3d visual grounding. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 19231–19242, 2023.
  • [48] Runsen Xu, Xiaolong Wang, Tai Wang, Yilun Chen, Jiangmiao Pang, and Dahua Lin. Pointllm: Empowering large language models to understand point clouds. In European Conference on Computer Vision, pages 131–147. Springer, 2024.
  • [49] Jintang Xue, Yun-Cheng Wang, Chengwei Wei, Xiaofeng Liu, Jonghye Woo, C-C Jay Kuo, et al. Bias and fairness in chatbots: An overview. APSIPA Transactions on Signal and Information Processing, 13(2), 2024.
  • [50] Le Xue, Mingfei Gao, Chen Xing, Roberto Martín-Martín, Jiajun Wu, Caiming Xiong, Ran Xu, Juan Carlos Niebles, and Silvio Savarese. Ulip: Learning a unified representation of language, images, and point clouds for 3d understanding. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pages 1179–1189, 2023.
  • [51] Zhengyuan Yang, Songyang Zhang, Liwei Wang, and Jiebo Luo. Sat: 2d semantics assisted training for 3d visual grounding. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 1856–1866, 2021.
  • [52] Zhenfei Yin, Jiong Wang, Jianjian Cao, Zhelun Shi, Dingning Liu, Mukai Li, Xiaoshui Huang, Zhiyong Wang, Lu Sheng, Lei Bai, et al. Lamm: Language-assisted multi-modal instruction-tuning dataset, framework, and benchmark. Advances in Neural Information Processing Systems, 36:26650–26685, 2023.
  • [53] Ting Yu, Xiaojun Lin, Shuhui Wang, Weiguo Sheng, Qingming Huang, and Jun Yu. A comprehensive survey of 3d dense captioning: Localizing and describing objects in 3d scenes. IEEE Transactions on Circuits and Systems for Video Technology, 34(3):1322–1338, 2023.
  • [54] Zhihao Yuan, Jinke Ren, Chun-Mei Feng, Hengshuang Zhao, Shuguang Cui, and Zhen Li. Visual programming for zero-shot open-vocabulary 3d visual grounding. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pages 20623–20633, 2024.
  • [55] Zhihao Yuan, Xu Yan, Yinghong Liao, Ruimao Zhang, Sheng Wang, Zhen Li, and Shuguang Cui. Instancerefer: Cooperative holistic understanding for visual grounding on point clouds through instance multi-level contextual referring. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 1791–1800, 2021.
  • [56] Tatiana Zemskova and Dmitry Yudin. 3dgraphllm: Combining semantic graphs and large language models for 3d scene understanding. arXiv preprint arXiv:2412.18450, 2024.
  • [57] Jirong Zha, Yuxuan Fan, Xiao Yang, Chen Gao, and Xinlei Chen. How to enable llm with 3d capacity? a survey of spatial reasoning in llm. arXiv preprint arXiv:2504.05786, 2025.
  • [58] Yiming Zhang, ZeMing Gong, and Angel X Chang. Multi3drefer: Grounding text description to multiple 3d objects. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 15225–15236, 2023.
  • [59] Lichen Zhao, Daigang Cai, Lu Sheng, and Dong Xu. 3dvg-transformer: Relation modeling for visual grounding on point clouds. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 2928–2937, 2021.
  • [60] Wayne Xin Zhao, Kun Zhou, Junyi Li, Tianyi Tang, Xiaolei Wang, Yupeng Hou, Yingqian Min, Beichen Zhang, Junjie Zhang, Zican Dong, et al. A survey of large language models. arXiv preprint arXiv:2303.18223, 1(2), 2023.
  • [61] Junsheng Zhou, Jinsheng Wang, Baorui Ma, Yu-Shen Liu, Tiejun Huang, and Xinlong Wang. Uni3d: Exploring unified 3d representation at scale. arXiv preprint arXiv:2310.06773, 2023.
  • [62] Ziyu Zhu, Xiaojian Ma, Yixin Chen, Zhidong Deng, Siyuan Huang, and Qing Li. 3d-vista: Pre-trained transformer for 3d vision and text alignment. In Proceedings of the IEEE/CVF International Conference on Computer Vision, pages 2911–2921, 2023.
  • [63] Ziyu Zhu, Zhuofan Zhang, Xiaojian Ma, Xuesong Niu, Yixin Chen, Baoxiong Jia, Zhidong Deng, Siyuan Huang, and Qing Li. Unifying 3d vision-language understanding via promptable queries. In European Conference on Computer Vision, pages 188–206. Springer, 2024.
@@ Appendix A Examples of Generated Descriptions
Figure 4: Qualitative Examples of Object-level Relational Descriptions Generated from Multi-view Images. The upper left part displays the image with object names, and the lower left part shows the bounding boxes of the objects.
We present qualitative examples of our object-level textual descriptions in Figure 4. Starting from detected object proposals and their corresponding multi-view images, we overlay the object names at the center of the projection areas in the image, as illustrated in the upper left part of the example. We then generate relational descriptions using a vision-language model. Key objects, typically those centrally positioned in the scene, are selected as query anchors. For each key object, we prompt the model to describe its spatial relationships with all other detected objects, resulting in detailed, contextually grounded descriptions. The prompt used is: “Describe clearly and briefly the relationships between the <Key Object>in the scene and nearby objects (<Other Object 1>, <Other Object 2>, …, <Other Object n>). Do not describe objects you cannot see.” For example, the objects in the image are a desk, two curtains, a window, a cabinet, and a table. There are two curtains, but only the one on the right is considered a key object because the other is positioned at the edge of the image. The chosen curtain is described as covering the window and situated near the table, the cabinet, and the desk. These relational descriptions offer interpretable summaries of local neighborhoods and equip downstream models with structured scene understanding for improved reasoning.
@@ Appendix B Ablation Study on Object Labels in Description Generation
To examine the impact of explicitly overlaying object category names during relational description generation, we conduct an ablation study comparing two variants: one where multi-view images include projected object labels (ours), and one without. As shown in Table 6, incorporating object labels consistently improves performance across all five benchmarks. The improvement is particularly notable in Scan2Cap and SQA3D, where more precise object references in the descriptions likely benefit caption generation and question answering. These results confirm that providing explicit category labels helps the vision-language model better ground each object and generate more informative relational descriptions.
@@
Appendix C Additional Quantitative Results

### v2

@@ 5 Conclusion
6 Acknowledgments
This project was sponsored by the US DoD LUCI (Laboratory University Collaboration Initiative) fellowship and the US Army Research Laboratory. The authors also acknowledge the Center for Advanced Research Computing (CARC) at the University of Southern California for providing computing resources.
@@ Appendix A Examples of Generated Descriptions
Figure 4: Qualitative examples of object-level relational descriptions generated using Prompt A (Default) with LLaVA-1.5
. The upper left part displays the image with object names, and the lower left shows the 3D bounding boxes. The right side contains relational descriptions emphasizing general proximity and object co-occurrence without enforcing strict spatial language.
Figure 5: Qualitative examples of object-level relational descriptions generated using Prompt B (Spatially Focused) with LLaVA-1.5. Compared to Prompt A, these descriptions include more explicit spatial terms (e.g., “on the left,” “behind”) and visual attributes, resulting in shorter but more positionally grounded sentences.
We present qualitative examples of our object-level textual descriptions in  fig.˜4. Starting from detected object proposals and their corresponding multi-view images, we overlay the object names at the center of the projection areas in the image, as illustrated in the upper left part of the example. We then generate relational descriptions using a vision-language model. Key objects, typically those centrally positioned in the scene, are selected as query anchors. For each key object, we prompt the model to describe its spatial relationships with all other detected objects, resulting in detailed, contextually grounded descriptions. The prompt used is: “Describe clearly and briefly the relationships between the <Key Object>in the scene and nearby objects (<Other Object 1>, <Other Object 2>, …, <Other Object n>). Do not describe objects you cannot see.” For example, the objects in the image are a desk, two curtains, a window, a cabinet, and a table. There are two curtains, but only the one on the right is considered a key object because the other is positioned at the edge of the image. The chosen curtain is described as covering the window and situated near the table, the cabinet, and the desk. These relational descriptions offer interpretable summaries of local neighborhoods and equip downstream models with structured scene understanding for improved reasoning.
@@ Appendix B Ablation Study on Object Labels in Description Generation
To examine the impact of explicitly overlaying object category names during relational description generation, we conduct an ablation study comparing two variants: one where multi-view images include projected object labels (ours), and one without. As shown in table˜7, incorporating object labels consistently improves performance across all five benchmarks. The improvement is particularly notable in Scan2Cap and SQA3D, where more precise object references in the descriptions likely benefit caption generation and question answering. These results confirm that providing explicit category labels helps the vision-language model better ground each object and generate more informative relational descriptions.
@@
Appendix C Ablation Study on Prompt Design for Object-Level Descriptions
To assess how different prompt formulations influence the quality of generated object-level relational descriptions and downstream 3D scene understanding, we compare two designs: a default prompt (Prompt A) that emphasizes relational conciseness, and a spatially grounded prompt (Prompt B) that encourages explicit spatial terms and appearance details.
fig.˜4 presents qualitative examples generated using Prompt A (Default). This prompt directs the vision-language model (LLaVA-1.5) to describe relational context in a concise, human-like way, without placing heavy emphasis on precise spatial markers. As shown, descriptions tend to mention object co-occurrence and proximity in natural, readable sentences. For example, “The curtain is covering the window, and it is also close to a table, a cabinet, and a desk.” While this phrasing lacks precise positional anchoring, it aligns with how humans intuitively describe contextual relevance.
In contrast, fig.˜5 illustrates the results of Prompt B (Spatially Focused). This prompt explicitly encourages the use of geometric relations (“on the left,” “in front of,” “behind”) and appearance details (“white,” “rectangular”), resulting in descriptions that are shorter but more spatially grounded. For instance, “The desk is located in a corner of the room… the window is above the desk… the cabinet is in front of the desk,” offers a clearer positional context but a less nuanced interpretation of function or co-usage.
To better understand how different prompts influence downstream model performance, table˜8 provides the full text of each design, and table˜9 summarizes quantitative results across five 3D vision-language tasks. While Prompt B encourages explicit spatial expressions (e.g., “to the left of,” “in front of”) and produces shorter sentences, it tends to focus narrowly on positional details, omitting functional or contextual cues. This results in descriptions that are more rigid but less informative overall. In contrast, Prompt A (Ours) generates a richer relational context with broader object co-occurrence and usage clues. As shown, Prompt A consistently outperforms Prompt B, suggesting that general, semantically rich descriptions better support multimodal reasoning than strictly spatial ones.
Prompt Version Prompt Text
Prompt A (Default) Describe clearly and briefly the relationships between the <key_object> in the scene and nearby objects (<other_obj1>, <other_obj2>, ...). Do not describe objects you cannot see. Do not describe green labels.
Prompt B (Spatially Focused) Based on the image, describe both the appearance and spatial relationships of the <key_object> in relation to nearby visible objects (<other_obj1>, <other_obj2>, ...). Include visual details like color, shape, size, or texture of the <key_object>, and explain precisely how it is positioned relative to nearby visible objects (<other_obj1>, <other_obj2>, ...) using terms such as ’on the left’, ’next to’, ’under’, ’in front of’, ’behind’, or ’on top of’. Only refer to what is clearly visible. Do not mention green text labels or objects not shown in the image."
Table 8: Comparison of prompt designs used for generating object-level relational descriptions with LLaVA-1.5. Prompt A is our default, concise formulation emphasizing relational grounding. Prompt B explicitly encourages spatial terms (e.g., “left,” “in front of”) and detailed appearance cues.
Prompt Design ScanRefer Multi3DRefer Scan2Cap ScanQA SQA3D
Acc@0.5 F1@0.5 C@0.5 CIDEr EM
Prompt B 51.4 55.1 74.1 92.3 55.2
Prompt A (Ours) 51.8 55.1 77.2 93.7 55.7
Table 9: Downstream performance using different prompts for generating object-level descriptions. Prompt B emphasizes spatial precision, while Prompt A (ours) encourages concise, general relational reasoning. Despite lacking explicit directional terms, Prompt A outperforms or matches Prompt B, suggesting that overly specific spatial descriptions may omit broader contextual signals useful for multimodal understanding.
Appendix D Ablation Study on Description Generator Choice
We further analyze the effect of the model used to generate object-level descriptions. While our main experiments adopt Vicuna-7B as the trainable backbone (to ensure fair comparison with prior methods such as Chat-Scene and 3DGraphLLM), the relational text can, in principle, be generated by any frozen captioner. table˜10 compares two options: generating descriptions with Vicuna-7B versus with LLaVA-1.5. Results show that replacing Vicuna with LLaVA-1.5 as the description generator improves downstream performance, especially on language-intensive tasks (e.g., +2.0 CIDEr on Scan2Cap, +2.1 CIDEr on ScanQA). This suggests that our framework is flexible with respect to the choice of description generator, and benefits from relation-dense captions produced by stronger multimodal models.
VLM ScanRefer Multi3DRefer Scan2Cap ScanQA SQA3D
Acc@0.5 F1@0.5 C@0.5 CIDEr EM
Vicuna-7B 51.8 55.0 75.2 91.6 55.6
LLaVa-1.5 (Ours) 51.8 55.1 77.2 93.7 55.7
Table 10: Impact of the choice of description generator. We compare object-level relational text produced by Vicuna-7B and by LLaVA-1.5 (used as frozen captioners). Using LLaVA-1.5 yields stronger downstream results.
Appendix E Additional Qualitative Results and Failure Cases
Additional Qualitative Examples
To further demonstrate the strengths of Descrip3D, we present additional qualitative comparisons of both question answering and object grounding tasks in fig.˜6. In the QA task fig.˜6(a), Descrip3D produces accurate answers in cases where Chat-Scene fails due to limited spatial awareness or insufficient contextual cues. For example, in the first question, while Chat-Scene incorrectly places the "single seat sofa" behind the brown chair, Descrip3D correctly identifies it as “in the corner of the room,” grounded by relational language. Similarly, Descrip3D succeeds in localizing queried objects such as the laptop and chairs based on complex object-to-object references, demonstrating its enhanced relational understanding. In the grounding task fig.˜6(b), Descrip3D resolves ambiguous references more reliably. For example, given a query like “the black couch next to a tall shelf and a fan,” Descrip3D identifies the correct object using spatial and contextual signals provided by the object descriptions. These results emphasize Descrip3D’s ability to perform robust reasoning in cluttered indoor environments where visual and geometric cues alone may be insufficient.
Failure Case Analysis
Despite its improved performance, Descrip3D is not immune to errors. fig.˜7 illustrates several representative failure cases in both QA (fig.˜7(a)) and grounding (fig.˜7(b)). In question answering, a common failure mode arises in counting tasks where performance is limited by upstream 3D detection accuracy. For instance, when the detector undercounts chairs around a table, Descrip3D cannot recover the correct answer solely through textual reasoning. Additionally, discrepancies between query phrasing (e.g., “square table”) and detector-generated object names (e.g., “coffee table”) introduce challenges in aligning language inputs with the available descriptions. In grounding, failures often occur when key attributes mentioned in the query (e.g., color or material) are missing or omitted in the generated object descriptions. As shown in  fig.˜7(b), the system fails to ground “the black cotton pillow” because the corresponding object description lacks explicit mention of its color, resulting in ambiguity during matching. These cases highlight limitations in both object detection accuracy and object description completeness, pointing to future directions for improving the robustness and coverage of relational grounding in 3D scene understanding.
(a) Additional qualitative comparison of question answering. Descrip3D correctly answers more challenging questions by leveraging precise spatial relations and context-aware relational cues from its textual descriptions.
(b) Additional qualitative comparison of grounding. Descrip3D provides more accurate object localization by resolving ambiguous object references using context-enhanced descriptions, such as "left of the fan" or “next to another door.”
Figure 6: Additional Qualitative comparison of 3D scene understanding tasks. Descrip3D outperforms Chat-Scene, especially in cases involving complex spatial grounding or multi-object reasoning, due to its use of a dual-level integrated relational textual descriptions that enhance contextual understanding.
(a) Additional qualitative comparison of question answering. Descrip3D fails in cases where the 3D detector misses objects or the query uses object names not aligned with detector output, limiting the effectiveness of textual reasoning.
(b) Additional qualitative comparison of grounding. Failures in grounding are often caused by incomplete object descriptions that omit key attributes (e.g., “black cotton pillow”), which are essential for accurate reference resolution.
Figure 7: Failure cases of 3D scene understanding tasks. While Descrip3D improves contextual reasoning, failure can still occur due to missing or ambiguous descriptions (e.g., color not mentioned) or mismatches between detection names and query expressions, especially in counting or spatial referencing.
Appendix F Additional Quantitative Results